OpenStudy (anonymous):

I have this trig question: If sec (theta) = -2 with and cos (theta) > 0. Find the six trigonometric functions of (theta) X=? Y=? r=? Sin (theta) = y/r

OpenStudy (anonymous):

csc (theta) = r/y cos (theta) = x/r Sec (theta) = r/x tan (theta) = y/x cot (theta) = x/y But I'm guessing that r=2 and x=-1 but I'm not sure

OpenStudy (anonymous):

I just have a problem what is X,Y and r. It's okay not to answer the 6 functions

OpenStudy (anonymous):

\(\sec\theta=\dfrac{1}{\cos\theta}=-2\), or \(\cos\theta=-\dfrac{1}{2}\). \(\cos\theta=\dfrac{x}{r}\), and since \(r\) is always positive, you have \(x=-1\) and \(r=2\). From this you can extract the value of \(y\), since \(x^2+y^2=r^2\). You'll then have all the info need to find the trig ratios. The information you're given, that \(\cos\theta>0\), doesn't make sense, since the cosine is clearly negative.